Image Segmentation by Refinement Reference manual

Version 1.2

Hend Ben Ameur ¡hbenameur@yahoo.ca¿Francois Clement ¡Francois.Clement@inria.fr¿Pierre Weis ¡Pierre.Weis@inria.fr¿

2017-11-23

1 Installation of Ref-image

To install the Ref-image command, follow the instructions in the INSTALL file of the distribution, or use the generic instructions of the next section.

Ref-image is free software distributed under the 3-clause BSD license, see the LICENSE file for details.

2 For the users of the distributed version of your package

They should follow the directives of the INSTALL file of the distribution of your package.

2.1 To configure, build and install the package

  ./configure [<installation root directory>]
  make
  make install

When called with one argument, the configure script makes it the root directory for the installation. When called without argument, the configure script interactively asks the user.

The variable installation root directory stands for the directory where the project will be installed; it must contain a valid directory name where the user can write. All the files that are installed by your project will be copied into sub directories of this one. Namely:

• binaries go into the sub-directory

     installation root directory/bin,

• includes go into the sub-directory

installation root directory/include/<project package name>,

• OCaml libraries go into the sub-directory

installation root directory/include/lib/ocaml/<project package name>

and other libraries go into the sub-directory

installation root directory/lib/<project package name>,

• manual pages go into the sub-directory

installation root directory/man/man<man section>,

• (La)TeX documentations go into the sub-directory

installation root directory/share/doc/tex and HTML documentation pages

go into the sub-directory

installation root directory/share/doc/htm,

• other materials go into the sub-directory

installation root directory/share/<project package name>.

2.2 To clean and remove the package

To remove all files installed by make install, all files built by make and all files created by ./configure, respectively type in:

make uninstall
make clean
./unconfigure